FEL Upgrade Project Weekly Brief
December 17-21, 2001



Highlights:
We wish all our friends and colleagues our best wishes for the holidays and the New Year.
We thank all who have helped us make the FEL Upgrade project and the FEL operations very successful over
the last year.  Jefferson Lab will shut down tonight for its annual holiday break and will re-open for business on
January 2nd.

This has been a very busy and very good week for us.  We placed a major procurement this week:  for the two
large dipole magnets that make up the major components in the recirculation bend.  We are pleased with both
the winning bidder and the contracted price.

We are pleased to announce that DODís Joint technology Office awarded two competitive grants to the FEL
project on Dec.18th: (1) the development of a high power drive laser for our advanced injectors ($280k in
FY02 with the option for a $570k follow-up ); and (2)the development of cryogenic mirrors for high power
FELs ($260k in FY02  with the option for $670k follow-up). Congratulations to George Neil, Steve Benson and
Michelle Shinn, who submitted the proposals.

We were also pleased to find out on Dec. 20th that the FY2002 Defense Appropriations Bill includes $1.7M to
continue the design and construction of the kilowatt class UV FEL which will be added to the 10 kW IR FEL
Upgrade.  This funding will come through our AFRL program office and will be shared with our partner on the
UV FEL project-Aerospace Corporation.  The appropriated funding is significantly less than the request
($5.5M) but it does keep the project moving forward during a period of tight budget constraints and we are
grateful to all who helped with the funding effort.

Management:
Project management was very busy with all that is noted in the highlights.  In addition we reviewed the end of year finances for the preparation of the project financial reports for November and December.

A draft version a technical proposal for a 100 kW Upgrade to the FEL facility was written and reviewed by the proposal team. prior to internal review.  This document has been requested by ONR and NAVSEA.

WBS 3 (Beam Physics):
Revision 1.1.3UV system design was completed and sent for layout & checking.  It features a UV bypass
design; this drove rearrangement of the linac-to-recirculator telescope so as to suppress geometric aberrations
in the transport to UV wiggler.  The reinjection telescope was tweaked to meet the match to the recirculator,
allowing for use of common designs.

GU/GV designs were reviewed and found to be acceptable in performance.  Documention will be generated
after the holiday closure.

The Design group met to discuss general progress.  PARMELA simulations in the hands of C. Hernandez,
activity on CSR under way, UV bypass design complete, and a very interesting presentation by S. Benson on
wiggler operation.

WBS 4 (Injector):
Drawings for the stalk rear roller guide and cross were signed off and submitted to the shop for fabrication.
The anode plate spider was machined.  We supplied a 3-inch diameter tube to W&M for a test implantation to
determine uniformity of material prior to implantation of the cathode support tube. Initial testing of the cathode
shield assembly showed very good non-binding movement in a vacuum environment

We held an internal design review of the preliminary design of the light box assembly supplied by Duly.  The
basic design looks good.  We made some simplifications for the next revision.

Gun HVPS - A vendor for the resistive inks, metalized terminations, and protective coatings has been found.
The question now is how to apply the various materials to the large ceramic cylinder to make the conditioning
resistor.  We also need to determine if the material can withstand a working gradient of 25-30 kV/inch.  We
will order some sample materials for testing.

WBS (SRF):
Received RF Power waveguide extensions.

WBS 6 (RF):
Quarter Klystrons - The first 100 kW Klystron for the upgrade was tested to 30 kW.  All parameters looked
great.  RadCon found a little radiation so the lead shielding is being reinstalled on the cart for ALARA.  The
unit will be run up to 50 kW and rechecked today.

Quarter HVPS - The design is complete and the parts are on order.  Some of the resistors will take 10-12
weeks to receive, but the construction can proceed without them.  The rigging company, Grubb, has not yet
told us how they will perform the lift to remove and then install the heavy components in the HVPS.  We will
get approval from M. Nevarez before proceeding with these lifts.

WBS 8 (Instrumentation):
This week began the real HARD work of the rip-out.  The cable trays in both the old and new optical cavity
areas have to be reworked.  This involves removing the 535MCM cable which powered up the optical chicane
and for the inject and extract strings. The 24 inch wide tray will be completely removed since it is no longer
needed on the linac side.  The jog in the tray around the collimator has to be removed to allow for the
waveguide for modules two and three. With a reduction of ~5 to 10 times in the number of trim magnets on the
linac side of the machine we will also remove all of those cables, as well as the optical diagnostic cables.  A
short video of the rip-out progress can be seen at http://laser.jlab.org and follow the "News" link or look at
http://laser.jlab.org/video/upgrade.

The new layout for stepper motors has been finalized.  There will be a total of 56 motors on the IR optical
cavity and the transport system with an additional 24 throughout the building for a total of 80 channels.  The
additional Oregon Micro Systems 8 channel VME board have been ordered.  The motor driver chassis are
being reworked from 4 channel to 8 channel.

The new shielded beam viewers are being prepared for movement to the vacuum assembly area in the test
lab.  This is where they will be installed as complete girder assemblies.  The flags are also being prepared with
6 micron Al foil, these will be "stamped" with grid marks at center, and + 5 & 10 mm in both the horizontal and
vertical.

Work on the new component data base continues.  All zones now have their designations and as components
are added or changed will be easily updated via the web site. This effort is essential to relocate the many
components that will be reused and will also serve as the basis of the new cable database.

WBS 9 (Transport):
Dipoles
Optical Chicane Dipoles (GW)
o  Magnet Enterprises International in Oakland CA. Potted the first article coil for the GW magnet
    and were able to demonstrate that the cure notice was effective and full production could begin
    again.
Injector Dipoles (DU/DV)
o The drawings and specifications were signed and Procurement is poised to go out for bid
    electronically in the first week of the year.
Arc Dipoles (GY, GX, GQ)
o Advanced Energy Systems (AES) is revised the drawings of the pole tips of the Bend Dipole
    (GQ) to reflect successful magnetic model results. All drawings are now complete.  David
    Douglas is reviewing the magnetic model results.  We will still want to analyze the model for
    reduced momentum at 145 and 80 MeV/c but will not let that delay the procurement process.
o We placed the contracts for the 180° Dipoles (DY) with Bosma Inc. of Ohio for the Cores and
    parts and Wang NMR of CA for coils and assembly.  The price was favorable to all.
Magnet Measurement
o Measurement of the corrector dipoles continued.
Quadrupoles
3 inch quad (QX)
o  The core vender, New England Techni-Coil (NTC) will use coils from the second batch to
    make 10 magnets which should be hear in the first days of the new year.
o  Jeffrey Hubrig of Innovation Services continued monitoring the production of QX Coils at
    Magnet Enterprises International (MEI).  The first batch under his advice came out well.  We
    have authorized MEI to continue full production.
Trim Quad (QT)
o  The magnets are in fabrication at Milhous Control of Virginia.
Sextupole (SF)
o  DULY Research continues to model the sextupole.  There were some questions about powering
    that were resolved and good results were seen in the 3D model by altering the exponent for the
    pole tip.  They were also authorized to decrease the size of the magnet to fit just over the beam
    pipe but still maintain good field over a 24 cm horizontal region.
Octupole (OT)
o  Work on this magnet is on furlough until the sextupole is designed.
Corrector Dipoles (DB, DJ)
o  The magnets continue to be qualified on the measurement stand.
Beam Line and Vacuum
o  Work is going very well at AES on the design the Arc Chambers.
o  Design work on girders and stands continued.

WBS 10 (Wiggler):
This week was involved in analyzing the commissioning data from wiggler 1.  We now have a downtime until
late January while the new vacuum chambers are tested.  After that the commissioning data will be taken for
wiggler 2 and then the final scans on the dispersion section.  To sum up the status at this point:

The wiggler vacuum chambers are done.  The supports for the diagnostics still need to be installed on the
wiggler support tube.  The commissioning data from wiggler 1 indicates that the wiggler meets its field integral
specifications over the complete range of excitation.  The kicks at the ends are quite small as well.  We still
need to check whether the fringe fields on the inner ends of the wigglers is affected by the presence of the
dispersion section.  This possible interference will have to be compensated by the dispersion section corrector
if it occurs.  Analysis of the spectra calculated from the entire wiggler, put together from the scans for each of
the three sections, indicate that one can create the exact spectrum of an ideal 25.5 period wiggler by operating
the dispersion section to produce minus 2.5 periods of net dispersion between the wigglers.  Operation for
production of X-rays also looks pretty straightforward.  Continuous tunability of the wavelength over a large
range looks possible if the dispersion section and wigglers are adjusted to keep the net dispersion in the
dispersion region constant.

WBS 11 (Optics):
Progress for the IR and UV upgrades were made in the following areas:

The drawing package for components from the "big C" (the yaw axis yoke) through the mirror holders is now
in review.  Based on modal analyses of the yaw plate, we are going to make the tensioning spring have a
variable preload, so we can tune the resonance frequency of the system.  A table giving the numbers and
specifications of all stepper motors used for the optical subsystems was generated for I&C.  One of the rotary
feedthroughs with the attached stepper motor (for the optical cavity assembly y-axis motion) was installed on a
short length of tubing to ensure UHV cleanliness, photographed and given to I&C for testing. A pair of ball
bearing assemblies were sent to Dicronite for coating.  We anticipate receiving the linear actuators and PZTs
for the optical cavity during the Winter shutdown.  Procurement has ordered the rails and pillow blocks for the
optical cavities (plus spares for the test stand).

The OCMMS beamline components have been redesigned and are ready for another review.  The OCMMS
tests have been moved to User Lab. 6, where we can take advantage of the hutch.  A fixture for the new
position sensitive detector was designed and sent to the Machine Shop for fabrication.

Discussions with one vendor on an advanced drive laser resulted in their being willing to provide amplifiers but
no oscillator.  Rather than "beat a dead horse", I have decided to drop the issue for now and concentrate on
another vendor.  From the available information, it should be straightforward to generate 40W of IR and
12-20W of green (depending on pulse width).  Talks continue and I anticipate receiving some budgetary
estimates during the shutdown.

Decommissioning/Other Activities:
The enclosure for the downstream optical table was removed, as were the HEPA filters for both optical
tables.  The Hall A Compton polarimeter team received some more assistance with the change out of the
mirrors in the high finesse cavity.  In return, we gained some valuable insight on some of the remote metrology
used, and its efficacy.  A brief report was generated as a result of that activity.  We continued data analyses
for the upcoming JTO review of FY01 activities